From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Feb 08 2003 - 16:11:29 EST
------------------------------------------------------------ Date: Sat Feb 8 16:11:26 2003, Shift: 08:00-16:00 Supervisor: ZM , with: GM Log book#14 pages 5 - 6 Shift Summary: ------------------------------------------------------------ 1. We got the beam for the entire shift. However from 11.00 the beam was very low and we voted to dump the beam togheter with PHENIX. Unfortunately MCR kept the beam due to tandem problems. The beam is dumped right now at 16:00. 2. Run 7838: acheved goal of 500k. Then we changed MRS polarity from A1050 to B1050. 3. No detector problems. NOTE: If collision rate is high or daq dead time is largen than 50% so set the scale down for trigger number 3 to two. ------------------------------------------------------------
This archive was generated by hypermail 2.1.5 : Sat Feb 08 2003 - 16:15:28 EST